# RUIE: Retrieval-based Unified Information Extraction using Large Language Model

### v1

@@ 4.2 Metrics
We employ span-based Micro-F1 to evaluate our method performance. For NER, an entity is considered correct if the entity span and type are correctly predicted. For RE, a relation is considered correct if relation type, subject entity, and object entity match the golden annotation. For EE task, we report two evaluation metrics: (1) ED: an event trigger is correct if the event type and the trigger word are correctly predicted. (2) EAE: an event argument is correct if its role type and event type match a reference argument mention.
@@ 4.3 Baseline Methods
We select three kinds of models for comparative analysis: (1) Traditional UIE method Lu et al. (2022), which is based on relatively small language models (i.e., million-level parameters). (2) Instruction-tuning based methods: Wang et al. (2023b) converts many public IE datasets to instruction-tuning datasets using natural language. Gui et al. (2024); Xiao et al. (2024b) constructs large instruction-tuning datasets with json output format. Li et al. (2024) uses code format to model IE structure. (3) Retrival-based methods: The first is sparse retriever bm25Robertson and Zaragoza (2009). The others are dense retriever Wang et al. (2024a); Xiao et al. (2024a) which are trained in a contrastive manner with large-scale text pair dataset. It is worth noting that Wang et al. (2023a); Wan et al. (2023) presuppose the entity span in the sentence, which is inconsistent with our experimental settings, so we did not compare them.
@@ 4.4 Implementation Details
Unless otherwise stated, we use llama3 as the scoring model and llama3.1-instruct for inference by default. For each test sample, we retrieve 8 examples for inference by default. We trained the keyword enhanced reward model and retriever based on ELECTRA-base and E5-base respectively. More detailed are in Appendix B:
@@ 5.1 Generalization to Held-out Tasks
In the held-out tasks generalization experiment, SFT-based methods are directly tested on held-out tasks after training on held-in tasks. After training on held-in tasks, RUIE retrieves demonstrations from held-out tasks for LLM. The remaining retrieval-based methods retrieve demonstrations for LLM directly at held-out tasks. Results are shown in Table 1.
RUIE achieves the best performance in four information extraction tasks, and demonstrates rapid generalization to new tasks. Compared with SFT-based methods, the generalization ability of RUIE is significantly better. The performance of SFT-based methods decreases seriously on datasets not seen during training, with the drop becoming more pronounced as task difficulty increases. For instance, in the ET and EAE tasks, the best-performing LLaMA2-IEPILE model only achieves F1 scores of 23.81 and 14.11, making it impractical for real-world use. In contrast, RUIE delivers improvements of 8.91, 14.89, 27.03, and 26.05 on NER, RE, ET, and EAE tasks, respectively, despite using smaller model.
Compared to general-purpose retrievers, RUIE retrieves higher-quality examples. BM25 is a strong baseline and outperforms semantic similarity based retrieval methods on average, confirming the alignment of fine-grained information is more important in IE example retrieval. By explicitly modeling fine-grained information, RUIE introduces LLMs preference as a supervision signal during retrieval training. Compared to BM25, which has the best overall performance, RUIE achieves 5.84, 4.05, and 3.07 improvements on NER, RE, and EET IE tasks respectively.
@@
Table 3: Performance (in F1%) comparison on RE tasks under held-in setting
@@
Table 4: Performance (in F1%) comparison on ED tasks under held-in setting.
@@
Table 5: Performance (in F1%) comparison on EAE tasks under held-in setting.
@@ 5.2 Results on Held-in Tasks
In the held-in tasks experiment, both supervised fine-tuning methods and RUIE are trained on held-in tasks and tested on their corresponding test sets. The remaining retrieving-based methods directly retrieve k examples from held-in tasks for LLM. SFT-based methods include Lu et al. (2022); Wang et al. (2023b); Gui et al. (2024); Xiao et al. (2024b); Li et al. (2024). As some baselines only report results for specific IE tasks, we report the SOTA results of the above methods in each dataset, denoted as “SOTA” in the tables.
The results of NER, RE, ED, and EAE are shown in Tables 2, 3, 4, and 5. We conclude two main findings: (1) SFT-basd methods provide LLMs with a comprehensive understanding of IE through large-scale IE datasets training, achieving superior results on held-in tasks compared to the RUIE, which utilizes a frozen LLM. However, RUIE achieved an F1 score of 92.26 on the FindVehicle, showing the strong potential of retrieval-based UIE. (2) RUIE showed markedly superior to generic retrievers, with improvements of 6.47, 7.13, 1.35, and 4.98 on the NER, RE, ED, and EAE tasks, respectively. This proves the effectiveness of introducing large model preference and keyword enhancement during the retriever training process.
@@ 5.3 Ablation Study
We conduct an ablation study across all held-out datasets to underscore the effectiveness of the key innovations in our work (Table 6). Removing the keyword enhancement led to a 0.62 decrease in the average F1-score across four tasks, highlighting the value of fine-grained information alignment during retriever training. After removing distill loss, the average f1-score of the four tasks decreases by 10.73, which is because the rank of candidate examples only reflects the relative distribution of LLs preferences. However, knowledge distillation based on KL-divergence can align the absolute distribution of LLMs preferences, which confirms the importance of using LLMs preferences as supervision signals. After removing the reward model, the average f1-score of the four tasks decreases by 8.87, indicating the log-likelihood from LLMs is not suitable as a direct knowledge source for distillation. Because the average log-likelihood is not a true probability distribution, and its values tend to cluster within a narrow range, making it less effective as a target distribution within the KL-divergence framework.
@@ 5.4 The Effects of Different K-shot Demonstrations
Figure 3: Performance (in F1%) comparison by varying k-shot demonstrations.
We use the default experimental setup to investigate the impact of varying k-shot demonstrations. Experiments were conducted across all held-in and held-out datasets, and we reported the average F1 scores for each task across all datasets, as depicted in Figure 3. We summarized three findings: (1) The performance across all tasks improved with an increase in k, with the most notable enhancement occurring as k rose from 0 to 2, indicating that few-shot demonstrations are crucial for LLMs to tackle tasks effectively. (2) The task performance does not consistently improve with the increment of k. For instance, there is a performance decline when k increased from 12 to 16 in the RE task. We surmise that one contributing factor is the limitation of the model’s context length, while another is the additional noise introduced by an excess of examples, which can adversely affect model performance. Therefore, we set k to 8, which is able to take into account both inference efficiency and extraction performance. (3) Model performance is influenced by the complexity of the task. NER has more data and is easier compared to other information extraction tasks, exhibited significantly better zero-shot performance than other tasks, while LLM was virtually incapable of completing the RE and ED tasks in a zero-shot scenario. Even as the number of examples k increased, the model’s performance on NER remained superior to that of other information extraction tasks.
@@ 5.5 The Effects of Different Scoring LLMs
Table 7: Performance (in F1%) comparison of different scoring LLMs on NER, RE, ED, EAE tasks across all held-out datasets.
To investigate the impact of various scoring LLMs on the extraction performance, we conducted experiments across all held-out datasets, with results presented in Table 7. We have two findings: 1) The size of the scoring LLM has a minor influence on the final performance. Although the LLaMA3 (8B) outperformed the GPTNeo (2.7B) by 2.98 in the ED task, their performance was comparable in NER, RE, and EAE tasks. This suggests that models of different sizes exhibit similar preferences for certain tasks, allowing one to select an appropriately sized LLM based on computational resources. 2) The base version of the LLM is more suitable as a scoring model than the instruct version. We calculated the mean and variance of the scores for the positive samples (top-3) and the negative samples (last-16) for both versions of the LLM (Table 10). Although the mean difference between positive and negative sample scores of instruct version LLM is larger, the variance of positive and negative sample scores is significantly larger than that of base version LLM, which introduces instability to the subsequent training of the reward model, resulting in the final extraction performance inferior to base version.
5.6 The Effects of Different Inference LLMs
Table 8 shows the performance of various reasoning LLMs across all held-out datasets. We find the ability of the reasoning LLMs significantly influence the extraction performance, which is expressed in two aspects: 1) Model size: For the same Qwen1.5, an increase in model size from 7B to 14B resulted in an average performance enhancement of 9.21 across four tasks. 2) Model type: Within the LLaMA series, LLaMA3.1, which is an advancement over LLaMA3, achieved an average performance improvement of 0.87 across four tasks. Since all the tasks are in English, the LLaMA series models consistently outperformed the Qwen series. Furthermore, RUIE can effectively adapt to both locally deployed models and API-based models, with Deepseek demonstrating the most superior performance.
@@ 5.6 The Effects of Different Inference LLMs
Table 8: Performance (in F1%) comparison of different inference LLMs on NER, RE, ED, EAE tasks across all held-out datasets. Deepseek refers to the deepseek-chat-v2. All other models are instruct variant, which are further fine-tuned on instruction-tuning datasets after pre-training.
@@ 6 Conclusion
In this paper, we propose RUIE, the first trainable retrieval-based framework designed for the UIE, which reduces the computational cost of fine-tuning and enables fast generalization to unseen tasks. We collected 31 datasets to train RUIE and tested it on 8 unseen tasks. The experimental results show that the introduction of LLMs preference and keyword-enhanced reward in RUIE effectively improves the information extraction ability of LLMs.
@@
Context length limitation: RUIE currently focuses on sentence-level UIE. On the one hand, it is difficult to meet the retrieval requirements of long documents due to the length limitation of the retriever is 512. On the other hand, due to the length limitation of LLMs, k-shot cannot be increased in the case of long documents, which limits the performance of the model.
@@ Appendix B Implementation Details
We finished LLM scoring, reward and retriever training on two 3090 GPUs. In order to balance the efficiency and performance of the scoring process, we used bm25 to initialize 100 candidates for each sample in the training set. Due to the context length limitation of the LLMs, we retrieve 8 samples for each query and set the maximum input length to 1792 and the maximum generation length to 256. We performed inference using vllm Kwon et al. (2023) on a single 3090 GPU. In order to ensure the reproduction of the results, we use a greedy decoding strategy and set the temperature to 0.
@@ Appendix C Score Mean and Variance of different LLMs
Table 10: Score Mean and Variance of different LLMs. “base” indicates LLaMA3-8B and “instruct” indicates LLaMA3-instruct-8B. We reported score mean and variance in positives and negatives separately.
@@
Appendix D Instruction Format

### v2

@@ 4.2 Metrics
We employ span-based Micro-F1 to evaluate the performance of our method. For NER, an entity is considered correct if the entity span and type are correctly predicted. For RE, a relation is considered correct if relation type, subject entity, and object entity match the golden annotation. For EE task, we report two evaluation metrics: (1) ED: an event trigger is correct if the event type and the trigger word are correctly predicted. (2) EAE: an event argument is correct if its role type and event type match a reference argument mention.
@@ 4.3 Baseline Methods
We compare our approach with three categories of methods:
Small-PLM based methods: Lu et al. (2022) proposed a unified framework based on medium-sized language models with task-specific instructions and structured prediction.
Instruction-tuning methods: These approaches fine-tune large language models with task-specific instructions. Wang et al. (2023b) reformulated the IE tasks using natural language instructions. Gui et al. (2024) and Xiao et al. (2024b) developed large-scale instruction datasets with JSON-formatted outputs. Li et al. (2024) encoded IE structures through code-like formats.
Retrieval-based methods: We include both sparse and dense retrievers. For sparse retrieval, we use BM25 Robertson and Zaragoza (2009). For dense retrieval, we employ state-of-the-art embedding models E5 Wang et al. (2024a) and BGE Xiao et al. (2024a), which are trained through contrastive learning on large-scale text pairs. We exclude methods that require pre-identified entity spans Wang et al. (2023a); Wan et al. (2023) to ensure fair comparison under our end-to-end extraction setting.
@@ 4.4 Implementation Details
We employ LLaMA3-8B as the scoring model and LLaMA3.1-8B-instruct for inference in our experiments by default. For each test instance, we retrieve 8 demonstrations by default. The keyword-enhanced reward model and retriever are implemented based on ELECTRA-base and E5-base, respectively. Detailed hyperparameters and training configurations are provided in Appendix B.
@@ 5.1 Generalization to Held-out Tasks
We evaluate the generalization capability of different approaches on held-out tasks. The SFT-based methods are directly tested on held-out tasks after training on held-in tasks. RUIE first trains on held-in tasks, then retrieves demonstrations from the mixed candidate pool for LLM on held-out tasks. The remaining retrieval-based methods directly retrieve demonstrations from the mixed candidate pool for LLM on held-out tasks. The results are presented in Table 1.
RUIE achieves the best performance in four information extraction tasks, and demonstrates rapid generalization to new tasks. Compared with SFT-based methods, the generalization ability of RUIE is significantly better. The performance of SFT-based methods decreases seriously on datasets not seen during training, with the drop becoming more pronounced as task difficulty increases. For example, in the ET and EAE tasks, the best performing LLaMA2-IEPILE model only achieves F1 scores of 23.81 and 14.11, making it impractical for use in the real world. In contrast, RUIE delivers improvements of 8.91, 14.89, 27.03, and 26.05 on NER, RE, ET, and EAE tasks, respectively, despite using a smaller model.
RUIE retrieves higher-quality examples compared to general-purpose retrievers. BM25 is a strong baseline and outperforms semantic similarity-based retrieval methods on average, confirming that fine-grained information alignment is more important in IE example retrieval. By explicitly modeling fine-grained information, RUIE introduces LLMs preference as a supervision signal during retrieval training. Compared to BM25, which has the best overall performance, RUIE achieves 5.84, 4.05, and 3.07 improvements on the NER, RE, and EET tasks, respectively.
RUIE demonstrates superior performance compared to simpler single-task corpus setups. When using general retrievers to retrieve examples for NER, there is a risk of retrieving RE or EE examples. In contrast, RUIE effectively minimizes the likelihood of retrieving examples from unrelated tasks. We conducted an experiment using BM25 to search within the candidate pool of the same task. Despite this simpler setup, RUIE achieved improvements of 1.79, 4.18, and 1.72 on the NER, RE, and EET tasks, respectively, even in multi-task candidate pool.
@@
Table 3: Performance (in F1-score) comparison on RE tasks under held-in setting $\bigstar$ indicates Wang et al. (2023b), $\spadesuit$ indicates Gui et al. (2024), $\blacktriangle$ incicates Xiao et al. (2024b), $\clubsuit$ indicates Li et al. (2024).
@@
Table 4: Performance (in F1-score) comparison on ED tasks under held-in setting. $\bigstar$ indicates Wang et al. (2023b).
@@
Table 5: Performance (in F1-score) comparison on EAE tasks under held-in setting. $\bigstar$ indicates Wang et al. (2023b), $\blacktriangle$ incicates Xiao et al. (2024b).
@@ 5.2 Results on Held-in Tasks
In the held-in tasks experiment, both the supervised fine-tuning methods and RUIE are trained on held-in tasks and tested on their corresponding test sets. The remaining retrieval-based methods directly retrieve k examples from held-in tasks for LLM. SFT-based methods include Lu et al. (2022); Wang et al. (2023b); Gui et al. (2024); Xiao et al. (2024b); Li et al. (2024). As some baselines only report results for specific IE tasks, we report the SOTA results of the above methods in each dataset, denoted as “SOTA” in the tables.
The results of NER, RE, ED and EAE are shown in Tables 2, 3, 4, and 5. We conclude two main findings: (1) SFT-basd methods provide LLMs with a comprehensive understanding of IE through training in large-scale IE datasets, achieving superior results on held-in tasks compared to RUIE, which utilizes a frozen LLM. However, RUIE achieved an F1 score of 92.26 on FindVehicle, showing the strong potential of retrieval-based UIE. (2) RUIE showed markedly superior performance to generic retrievers, with improvements of 6.47, 7.13, 1.35, and 4.98 on the NER, RE, ED, and EAE tasks, respectively. This shows the effectiveness of introducing large model preference and keyword enhancement during the retriever training process.
We have also identified several key error patterns: 1. Domain knowledge gaps: Poor performance on medical datasets (e.g. bg2gm, SciERC) due to LLaMA3.1’s limited medical domain training. 2. Label confusion: Difficulty distinguishing similar target labels (e.g. FabNER, kbp37). 3. Text formatting: Decreased precision with massive special symbols such as “@” and “#” in HarveyNER. 4. Pronoun handling: Missed pronoun entities in ACE2005. Moreover, we observed that the information extracted by the LLM and the actual labels only differ by one or two words, yet this does not hinder human understanding of the extraction results. This suggests that LLM-based information extraction methods require a more nuanced evaluation metric.
@@ 5.3 Ablation Study
We conducted an ablation study across all held-out datasets to underscore the effectiveness of the key innovations in our work (Table 6). Removing the keyword enhancement led to a 0.62 decrease in the average F1-score across four tasks, highlighting the value of fine-grained information alignment during retriever training. After removing distill loss, the average F1-score of the four tasks decreases by 10.73, which is because the rank of candidate examples only reflects the relative distribution of LLMs preferences. However, knowledge distillation based on KL-divergence can align the absolute distribution of LLMs preferences, which confirms the importance of using LLMs preferences as supervision signals. After removing the reward model, the average F1-score of the four tasks decreases by 8.87, indicating the log-likelihood from LLMs is not suitable as a direct knowledge source for distillation. The reason is that the average log-likelihood is not a true probability distribution, and its values tend to cluster within a narrow range, making it less effective as a target distribution within the KL-divergence framework.
@@ 5.4 The Effects of Different K-shot Demonstrations
Figure 3: Performance (in F1-score) comparison by varying k-shot demonstrations.
We first used the default experimental setup to investigate the effects of varying k-shot demonstrations on extraction performance. Experiments were conducted across all held-in and held-out datasets, and we reported the average F1 scores for each task across all datasets, as depicted in Figure 3. We summarized three findings: (1) The performance across all tasks improved with an increase in k, with the most notable enhancement occurring as k rose from 0 to 2, indicating that few-shot demonstrations are crucial for LLMs to tackle tasks effectively. (2) The task performance does not consistently improve with increasing k. For example, there is a performance drop when k increased from 12 to 16 in the RE task. We surmise that one contributing factor is the limitation of the model’s context length, while another is the additional noise introduced by an excess of examples, which can adversely affect model performance. Therefore, we set k to 8, which can take into account both inference efficiency and extraction performance. (3) Model performance is influenced by the complexity of the task. NER has more data and is easier compared to other information extraction tasks, exhibited significantly better zero-shot performance than other tasks, while LLM was virtually incapable of completing the RE and ED tasks in a zero-shot scenario. Even as the number of examples k increased, the model’s performance on NER remained superior to that of other information extraction tasks.
Figure 4: Inference speed (instances per second) comparison of different k-shots demonstrations on ACE 2004 (NER), ADE corpus (RE) and ACE2005 (ED and EAE) datasets.
Then we used the default experimental setup to investigate the effects of different k-shots on inference efficiency. Since SFT-based methods with similar size do not support vLLM, we treat the case of k-shot = 1 as equivalent to the SFT-based methods for fairness (RUIE is built based on bi-encoder and the retrieval time of a single sample can be negligible in our experiment). As shown in Figure 4, the inference speed decreases consistently with increasing number of k shots, with k-shot = 1 approximately 1.9 times faster than k-shot = 8. While RUIE has lower inference efficiency than SFT-based due to retrieval overhead and longer context length, it offers three key advantages: 1. Minimal training cost: Only fine-tunes a small retriever vs. full LLM. 2. Better domain adaptation: Using ICL with a small amount of labeled data vs. full retraining. 3. API compatibility: Works directly with LLM APIs vs. requiring model deployment.
@@ 5.5 The Effects of Different Scoring LLMs
Table 7: Performance (in F1-score) comparison of different scoring LLMs on NER, RE, ED, EAE tasks across all held-out datasets.
To investigate the impact of various scoring LLMs on the extraction performance, we conducted experiments across all held-out datasets, with results presented in Table 7. We have two findings: 1) The size of the scoring LLM has a minor influence on the final performance. Although LLaMA3 (8B) outperformed GPTNeo (2.7B) by 2.98 on ED task, their performance was comparable on NER, RE and EAE tasks. This suggests that models of different sizes exhibit similar preferences for certain tasks, allowing one to select an appropriately sized LLM based on computational resources. 2) The base version of the LLM is more suitable as a scoring model than the instruct version. We calculated the mean and variance of the scores for the positive samples (top-3) and the negative samples (last-16) for both versions of the LLM (Table 10). Although the mean difference between positive and negative sample scores of instruct version LLM is larger, the variance of positive and negative sample scores is significantly larger than that of base version LLM, which introduces instability to the subsequent training of the reward model, resulting in the final extraction performance inferior to base version.
5.6 Performance Analysis Across Different LLMs for Inference
Table 8 shows the performance of various reasoning LLMs across all held-out datasets. The ability of the inference LLMs significantly influences the extraction performance, which is expressed in two aspects: 1) Model size: For the same Qwen1.5, an increase in model size from 7B to 14B resulted in an average performance enhancement of 9.21 across four tasks. 2) Model type: Within the LLaMA series, LLaMA3.1, which is an advancement over LLaMA3, achieved an average performance improvement of 0.87 across four tasks. Since all tasks are in English, the LLaMA series models consistently outperformed the Qwen series. Additionally, RUIE adapts effectively to both locally deployed models and API-based models, with Deepseek demonstrating the best performance overall.
@@ 5.6 Performance Analysis Across Different LLMs for Inference
Table 8: Performance (in F1-score) comparison of different inference LLMs on NER, RE, ED, EAE tasks across all held-out datasets. Deepseek refers to the deepseek-chat-v2. All other models are instruct variant, which are further fine-tuned on instruction-tuning datasets after pre-training.
@@ 6 Conclusion
In this paper, we introduced RUIE, a novel trainable retrieval-based framework for unified information extraction that addresses key challenges in the field. Our framework introduced a pioneering trainable retrieval mechanism specifically designed for UIE tasks, significantly reducing computational costs while enabling rapid generalization to unseen tasks. Extensive experiments on 31 training datasets and 8 held-out tasks demonstrated RUIE’s superior performance across various IE tasks, suggesting promising directions for developing more efficient and adaptable IE systems.
@@
Sequence length constraint: RUIE currently focuses on sentence-level UIE. On the one hand, it is difficult to meet the retrieval requirements of long documents due to the length limitation of the retriever is 512. On the other hand, due to the length limitation of LLMs, k-shot cannot be increased in the case of long documents, which limits the performance of the model.
@@
Acknowledgments
This work was supported in part by the National Key Research and Development Program of China (No.2021YFF1201200), the Science and Technology Major Project of Changsha (No.kh2402004). This work was carried out in part using computing resources at the High-Performance Computing Center of Central South University.
@@ Appendix B Implementation Details
We finished LLM scoring, reward and retriever training on two 3090 GPUs. In order to balance the efficiency and performance of the scoring process, we used bm25 to initialize 100 candidates for each sample in the training set. Due to the sequence length limitation of the LLMs, we retrieved 8 samples for each query and set the maximum input length to 1792 and the maximum generation length to 256. We performed inference using vLLM Kwon et al. (2023) on a single 3090 GPU. In order to ensure the reproduction of the results, we use a greedy decoding strategy and set the temperature to 0.
@@ Appendix C Score Mean and Variance of different LLMs
Table 10: Score Mean and Variance of different LLMs. “base” indicates LLaMA3-8B and “instruct” indicates LLaMA3-8B-instruct. We reported score mean and variance in positives and negatives separately.
@@
Appendix D Sample Format
Detailed sample format is listed in Table 12.
Appendix E Instruction Format
@@
Task Sample Format
NER Task: Named Entity Recognition
Schema: [location, person, organization]
Input: The Parkinsons are a punk rock band originally from Coimbra, Portugal, formed in the year 2000 and based in London, known for their outrageous live performances.
Output: location: Coimbra; location: Portugal; location: London.
RE Task: Relation Extraction
Schema: [Organization based in, Located in, Live in, Work for, Kill]
Input: Washington: About 110 firefighters cut a containment line most of the way around an 850-acre forest fire in the Pasayten Wilderness near the Canadian border Tuesday.
Output: Located in: Pasayten Wilderness, Washington.
ED Task: Event Detection
Schema: [phishing, data breach, ransom, discover vulnerability, patch vulnerability]
Input: Google Project Zero’s security researchers have discovered another critical remote code execution (RCE) vulnerability in Microsoft’s Windows operating system, claiming that it is something truly bad.
Output: discover vulnerability: have discovered.
EAE Task: Event argument extraction
Schema: Given event trigger: “discover vulnerability: discovered”; Candidate arguments: [vulnerable system owner, vulnerability, capabilities, time, vulnerable system version, discoverer, common vulnerabilities and exposures, supported platform, vulnerable system]
Input: Google Project Zero’s security researchers have discovered another critical remote code execution (RCE) vulnerability in Microsoft’s Windows operating system, claiming that it is something truly bad.
Output: vulnerable system: Windows operating system; vulnerability: remote code execution (RCE) vulnerability; discoverer: security researchers; vulnerable system owner: Microsoft; discoverer: Google Project Zero.
Table 12: Detailed sample format across four IE tasks. NER sample is from polyglotNER Al-Rfou et al. (2014), RE sample is from conll04 Roth and tau Yih (2004), and both ED and EAE samples are from CASIE Lu et al. (2021).
@@
2. Include Only Present Relations: Only output relations that actually exist in the input. Ignore any relations that are not mentioned.